OpenStudy (anonymous):

A = quantity of B plus K all over G A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = quantity of B plus K all over G for K. Please show your work. (6 points) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number. (4 points) Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

A=(B+K)/G AG=B+K) AG-B=K